While observing A085524 the lamp failed, sometimes there were too few counts and sometimes there were no counts (like a bias). When I tried to turn on the lamp manually in the GUI I received the error message:
GUI ERROR ON COMAND LAMP
set TH-AR OFF TIME-OUT on replay for command
LAMP set TH-AR OFF ==> /OPTGUI_Device.vi/
ENV_SendCommandWaitReply.vi:154/
ERR_AddErrorlfError.vi
After of "playing" with the GUI and the commands for about one hour I was able to take the lamps again, so I could continue with the observations… BUT, the problem came back several times during the night. Turning ON the lamp manually and then taking exposing it was possible to get the calibration lamps (not always), so as a partial solution I did it in order to continue with the observations, several spectra are named "junk" because of this problem.
The exposure meter shows different total counts for the lamp calibrations, I do not know if this issue is associated with the same problem.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-06-26 | 9.25 | 0.0 | 1.25 | 1.0 | 0.0 | 11.5 |
High humidity the first quarter of the night.
Please unmark A085524 from the NOS.
Targets not observed: HIP058099, A091704, HIP073258 and HIP08622.